# QLLM: Do We Really Need a Mixing Network for Credit Assignment in Multi-Agent Reinforcement Learning?

### v4

@@ Introduction
This section presents a detailed discussion of the implementation of each component of QLLM. We first propose the concept of the IGM-Gating Mechanism, which enables the learning architecture to switch between strict adherence to the IGM principle and more relaxed joint value modeling. This provides a foundation for adaptively handling different types of multi-agent tasks. Then we introduce the concept of TFCAF, a training-free credit assignment function capable of accurately assigning credit to individual agents. Leveraging the generative capabilities of LLMs, TFCAF is constructed through prompt-based interaction with LLMs. However, when applied to complex tasks, LLMs typically encounter two fundamental challenges, hallucination (Tonmoy et al. 2024) and a lack of rigorous reasoning. These limitations motivate the design of our proposed coder-evaluator framework. By integrating this framework, QLLM is able to generate highly accurate TFCAF functions for credit assignment in multi-agent settings, offering an effective alternative to the traditional mixing network used in value decomposition approaches.
IGM-Gating Mechanism
To flexibly handle diverse coordination demands across tasks, we introduce the IGM-Gating Mechanism, which enables the learning architecture to switch between strict adherence to the IGM principle and relaxed joint value modeling. This mechanism allows for conditional enforcement of the monotonicity constraint in the credit assignment.
Definition 1 (IGM-Gating Mechanism).
Let $Q_{\text{tot}}(\boldsymbol{s},\boldsymbol{a})$ be the joint action-value function, and $\{Q_{i}(s_{i},a_{i})\}_{i=1}^{n}$ the individual utility functions. Define a binary gating variable $\gamma\in\{0,1\}$, where $\gamma=1$ indicates that IGM is enforced. The IGM-Gating Mechanism imposes the following conditional monotonicity constraint:
$\forall i,\quad\gamma=1\Longrightarrow\frac{\partial Q_{\text{tot}}(\boldsymbol{s},\boldsymbol{a})}{\partial Q_{i}(s_{i},a_{i})}\geq 0.$ (3)
When the gating variable is inactive ($\gamma=0$), no monotonicity constraint is enforced, allowing $Q_{\text{tot}}$ to freely capture non-monotonic dependencies.
Training-Free Credit Assignment Function
As a crucial component of credit assignment, the mixing network in value decomposition is essentially a state-conditioned feedforward neural network composed of multiple linear transformation layers and nonlinear activation functions. The network takes as input the individual Q-values of all agents, denoted as $\boldsymbol{Q}=\{Q_{1},Q_{2},\dots,Q_{n}\}$, and produces the joint Q-value $Q_{\text{tot}}$ as output. To enable the network to be state-dependent and adaptively parameterized, most value decomposition methods employ the hypernetwork mechanism (Ha, Dai, and Le 2017). Specifically, instead of being directly optimized, the weights and biases of each linear layer in the mixing network are generated by a hypernetwork that conditions on the global state $s$. This design facilitates state-aware adaptation while ensuring compatibility with CTDE framework, making deployment straightforward. However, such neural-network-based credit assignment functions cannot effectively incorporate prior knowledge, which may result in slower convergence and reduced interpretability.
@@
To address these issues, we propose the Training-Free Credit Assignment Function (TFCAF), which is the core module of QLLM. This module is generated by LLMs through iterative error correction, based on a series of easily accessible task-related prompts. As illustrated in Equation (2), the global Q-value function $Q_{\text{tot}}$ can theoretically be decomposed into a weighted sum of individual local Q-values with an additional bias term. Motivated by this decomposition paradigm, we incorporate a similar formulation in the design of TFCAF, which we formally define below.
Definition 2.
A Training-Free Credit Assignment Function (TFCAF) is defined as:
$Q_{\text{tot}}\left(s,\boldsymbol{a}\right)=\sum_{i=1}^{n}f_{w}^{i}(s)\,Q_{i}\left(\tau^{i},a^{i}\right)+f_{b}(s),$ (4)
where $f_{w}^{i}(s)$ and $f_{b}(s)$ are general nonlinear functions of the global state $s$ generated by LLMs before training.
As illustrated in Definition 1, when the gating signal $\gamma=1$, the function $f_{w}^{i}(s)$ is strictly constrained to be non-negative. In contrast, when $\gamma=0$, this constraint is lifted. Benefiting from the rich prior knowledge of LLMs, the appropriate value of $\gamma$ can be easily inferred from the task description, enabling the model to adapt to different types of tasks. For instance, in matrix game scenarios, where individual optimal decisions do not necessarily align with the global optimum, the LLM may assign a negative value to $f_{w}^{i}(s)$. The coefficients $f_{w}^{i}\left(s\right)$ and the bias term $f_{b}\left(s\right)$ make the overall TFCAF a nonlinear mapping, although its final combination with the local $Q_{i}$ remains linear in form. Both components are automatically produced by LLMs through the coder-evaluator framework, which we detail next along with the iterative correction and its role in credit assignment for multi-agent systems.
Coder Evaluator Framework
Due to the susceptibility of LLMs to hallucinations, the process of generating TFCAF must be designed to be robust. To address this challenge, we propose the coder-evaluator framework, which automatically generates TFCAF through a modular and self-correcting process, ensuring both accurate credit assignment and logical consistency in code generation.
This framework employs two LLMs with distinct yet complementary roles: the coder LLM $M_{coder}$, responsible for generating candidate TFCAFs, and the evaluator LLM $M_{evaluator}$, which critically evaluates these candidates. All of these are generated by LLMs through structured prompts without any need for human intervention.
@@ A Training-Free Credit Assignment Function (TFCAF) is defined as:
The interaction between the two models is driven by two types of prompts: task prompts ($P_{task}$) specify the meanings of each dimension in the global state space, the action space of the agents, and the composition of the reward function. This information is objective and can be easily obtained from the environment; role prompts ($P_{role}$) define the fixed responsibilities and operational roles of each LLM within the framework. The role prompts are further categorized into coder prompts and evaluator prompts, which remain unchanged across different RL tasks and correspond to the Coder LLM and the Evaluator LLM, respectively. This clear separation and stability of prompt design make QLLM easy to adapt to new cooperative tasks without extensive manual prompt tuning or subjective rewriting (see in Appendix A).
@@
Initial Generation.
Firstly, $M_{coder}$ generates $K$ candidate credit assignment functions by processing both the role-specific and task-specific prompts:
$\phi_{1},\phi_{2},...,\phi_{K}\ \Leftarrow\ M_{coder}\big(P_{role},\ P_{task}\big),$ (5)
where $\phi_{1},\phi_{2},\dots,\phi_{K}$ denote candidate TFCAFs, each including instructions for calculating state-dependent weights and biases.
Error Detection and Correction.
Due to the inherent hallucination tendencies of LLMs, candidate functions might contain syntax or dimensional errors. To address this, each candidate TFCAF is compiled and run once with any available input state and local Q-values. If an error occurs during execution, the error message generated by the compiler will be fed back to $M_{coder}$, prompting regeneration of the faulty TFCAF:
$\phi_{correct}\,\,\Leftarrow\,\,M_{coder}\big(P_{role},\,\,P_{task},\ error\big).$ (6)
This correction step repeats until all $K$ candidate TFCAFs can execute successfully without errors.
Selection of TFCAF.
Subsequently, $M_{evaluator}$ evaluates these validated candidate TFCAFs, selecting the most promising one based on task relevance, interpretability, and code clarity. Notably, the selection of the best candidate is fully guided by structured prompts and performed automatically by $M_{evaluator}$, without any manual or subjective intervention:
$choice\,\Leftarrow\,\,M_{evaluator}\big(P_{role},\ P_{task},\ {\Phi}\big),$ (7)
where ${\Phi}=\{\phi_{1},\phi_{2},...,\phi_{K}\}$.
Algorithm 1 Pseudo-code for QLLM
1:Part 1: TFCAF Generation (LLM-based)
2: $\phi_{1},\phi_{2},\dots,\phi_{K}\leftarrow M_{\text{coder}}(P_{\text{role}},\,P_{\text{task}})$ $\triangleright$ Generate $K$ candidate functions
3:for all $\phi_{k}\in\{\phi_{1},\dots,\phi_{K}\}$ do
4:  if $\phi_{k}$ fails on environment samples then
5:   $\phi_{k}\leftarrow M_{\text{coder}}(P_{\text{role}},P_{\text{task}},\text{error})$ $\triangleright$ Regenerate function based on error message
6:  end if
7:end for
[…]
10:Generate $f_{\text{TFCAF}}$ from $\phi_{\text{final}}$ using predefined template
11:Part 2: Agents Optimization (RL-based)
12:for episode = 1 to $M$ do
13:  for each agent $i$ do
14:   $a^{i}\leftarrow\arg\max_{a}Q_{i}(\tau^{i},a;\theta)$
15:  end for
16:  Compute global Q-value:
17:   $Q_{\text{tot}}(s,\boldsymbol{a})=f_{\text{TFCAF}}(Q_{1},...,Q_{n};s)$
18:  Store transition $(s,\boldsymbol{a},r,s^{\prime})$ in buffer $\mathcal{D}$
19:  Sample batch and compute TD target:
20:   $y=r+\gamma\max_{\boldsymbol{a}^{\prime}}Q_{\text{tot}}(s^{\prime},\boldsymbol{a}^{\prime};\theta^{-})$
21:  Update parameters $\theta$ by minimizing the TD loss:
22:   $\mathcal{L}(\theta)=\mathbb{E}\left[\left(y-Q_{\text{tot}}(s,\boldsymbol{a};\theta)\right)^{2}\right].$
23:end for
Training Procedure
@@
where $f_{\text{TFCAF}}$ is a parameterized function in which the weights are allowed to take arbitrary values, while the biases are also unconstrained. $f_{\text{TFCAF}}$ receives the agent-specific Q-values $Q_{i}$ as input and utilizes the global state $s$ to generate the corresponding weights and biases through the weight module $f_{w}^{i}$ and the bias module $f_{b}$, respectively. Notably, all parameters in the TFCAF are generated by LLMs and remain fixed during training. The global Q-value is then computed as shown in Equation (4). The training objective is to minimize the temporal-difference (TD) loss between the predicted total Q-value and the target Q-value:
@@
where the target value $y$ is computed using Bellman equation:
@@
with $\gamma$ being the discount factor. $\theta^{-}$ denotes the parameters of the target agent network.
@@
Experiments
Figure 3: Average episodic return curves for selected tasks in GRF and MPE environments (the results for LBF are shown in Appendix C).
To comprehensively evaluate the proposed method, the experiments were designed with six main objectives: (1) to assess the performance of the QLLM algorithm in terms of average return across multi-agent environments; (2) to verify its adaptability to high-dimensional global state spaces by increasing the dimensionality in selected tasks; (3) to demonstrate the compatibility of QLLM by integrating it into other mixing-based algorithms; (4) to demonstrate the compatibility of QLLM with various LLMs; (5) to validate the effectiveness of the proposed coder-evaluator framework under various parameters; (6) to demonstrate the superiority of our method in terms of training costs.
Experimental Setups
@@
We tested QLLM’s performance on four common MARL benchmarks and compared it with several well-known baseline credit assignment algorithms. The primary LLM used in our method is the DeepSeek-R1 (Guo et al. 2025) inference model. For each experiment, the reported return values represent the mean over five independent training runs, with 90% confidence intervals indicated by error bars. Unless otherwise specified, all environment parameters are set to their default values. Our experimental environments include: (1) Cooperative Matrix Games [^1] https://github.com/uoe-agents/matrix-games, used to validate the algorithm’s performance in non-monotonic scenarios, which may violate the IGM principle; (2) Multi-agent Particle Environments (MPE) [^2] https://pettingzoo.farama.org/environments/mpe/, for evaluating performance in high-dimensional state spaces; (3) Level-Based Foraging (LBF) [^3] https://github.com/uoe-agents/lb-foraging, for testing in sparse reward settings; and (4) Google Research Football (GRF) [^4] https://github.com/google-research/football, to assess performance in complex, realistic environments.
@@
We compare QLLM with six commonly used baseline algorithms: QMIX (Rashid et al. 2020b), QPLEX (Wang et al. 2020), Weighted QMIX (Rashid et al. 2020a), Qatten (Yang et al. 2020), RIIT (Hu et al. 2021), MASER (Jeon et al. 2022), and COMA (Foerster et al. 2018). A detailed introduction and implementation details for all baselines and environments are provided in Appendix  B.
Figure 4: Average episodic return curves in matrix-games environment (Right) and the payoff matrix used in this environment (Left).
Figure 5: Compatibility evaluation of QLLM in the MPE environments. The original mixing networks in RIIT, MASER, and QMIX are replaced by the TFCAF generated by QLLM, and the resulting performance is compared against the original implementations.
The Superiority of QLLM
Average Return and Convergence Speed.
Figure 3 and Figure 4 show the performance of various algorithms in certain scenarios. It is evident that the QLLM algorithm outperforms the baseline in both the dense-reward MPE environment and the sparse-reward GRF environment. Especially in the GRF environment, COMA almost fail to train the agents properly, which highlights the importance of accurate credit assignment in multi-agent environments. Furthermore, the incorporation of the IGM-Gating mechanism enables our algorithm to perform effectively in non-monotonic matrix game environments, outperforming methods like QMIX that are limited by rigid monotonicity constraints. In Listing 1, we present an example of a credit assignment function generated by LLMs for the pz-mpe-simple-spread task in the MPE environment. It can be seen that this function takes into account key factors such as the agents’ positions, their distances to landmarks, as well as the number of collisions, accurately assigning contributions to each agent. These verify the effectiveness and generalization ability of the credit assignment functions generated by our method.
Figure 6: The results of high-dimensional testing in MPE environment.
Performance in Higher Dimensional State Spaces.
We tested the performance of various algorithms in high-dimensional state spaces of certain scenarios. Taking the pz-mpe-simple-spread task as an example in the MPE environment, in the navigation scenario, as the number of agents increases, the global state space dimension becomes very large, and the difficulty of credit assignment grows exponentially. In Figure 6, the results indicate that most algorithms perform poorly when faced with such a high-difficulty task, while our algorithm still maintains a high level of contribution assignment accuracy, which is attributed to the mechanism of using LLMs in QLLM. The baseline algorithms rely entirely on neural networks for credit assignment, and as the state dimensions increase, determining the weight values becomes more difficult. However, our algorithm uses a set of well-defined nonlinear functions to assign contributions, which are unaffected by high-dimensionality, thus maintaining a high level of performance.
Listing 1: A credit assignment function example of pz-mpe-simple-spread environment generated by LLMs.
def QLLMNetwork(agents_q: torch.Tensor, global_state: torch.Tensor) -> torch.Tensor:
batchsize = agents_q.size(0)
n_agents = 6
global_state_reshaped = global_state.view(batchsize, n_agents, 36)
# Process landmark distances
landmarks_relative = global_state_reshaped[:, :, 4:16].view(batchsize, n_agents, 6, 2)
landmark_distances = torch.norm(landmarks_relative, dim=-1)
min_dist_per_agent = torch.clamp(torch.min(landmark_distances, dim=-1)[0], min=0.1)
pos_weights = 1.0 / min_dist_per_agent
# Process collision counts
positions = global_state_reshaped[:, :, 2:4]
diffs = positions.unsqueeze(2) - positions.unsqueeze(1)
pairwise_dist = torch.norm(diffs, dim=-1)
collision_mask = (pairwise_dist <= 0.3) & ~torch.eye(n_agents, device=positions.device).bool().unsqueeze(0)
collision_weights = collision_mask.sum(dim=-1).float()
# Combine weights and softmax
combined_weights = pos_weights - collision_weights
softmax_weights = torch.softmax(combined_weights, dim=-1)
global_q = (agents_q * softmax_weights).sum(dim=-1, keepdim=True)
return global_q
Compatibility Test.
@@
QLLM is still based on value decomposition methods, so it is compatible with most MARL algorithms that use mixing networks to compute global Q-values. We used the TFCAF generated by QLLM to replace the mixing networks in three value-decomposition algorithms—RIIT, MASER, and QMIX. We then conducted experiments in the MPE environment to examine how integrating TFCAF affects the performance of these value-decomposition MARL methods. The experimental results tested in MPE environments are shown in Figure 5. Our results show that QLLM can be seamlessly incorporated into existing architectures, offering a flexible and scalable solution for a wide range of multi-agent settings.
Conclusion

### v5

@@ 4 Methodology
This section presents a detailed discussion of the implementation of each component of QLLM, whose overall framework is illustrated in Figure 2. We first introduce the concept of TFCAF, a training-free credit assignment function capable of accurately assigning credit to individual agents. Leveraging the generative capabilities of LLMs, TFCAF is constructed through prompt-based interaction with LLMs. However, when applied to complex tasks, LLMs typically encounter two fundamental challenges, hallucination [39] and a lack of rigorous reasoning. These limitations motivate the design of our proposed coder-evaluator framework. By integrating these components, QLLM is able to generate highly accurate TFCAF functions for multi-agent credit assignment, providing an effective alternative to the traditional mixing networks used in value decomposition approaches.
Figure 2: QLLM framework: $M_{\text{coder}}$ generates candidate TFCAFs, and $M_{\text{evaluator}}$ selects the optimal function after automated verification.
4.1 Training-Free Credit Assignment Function
Traditional mixing networks often rely on hypernetwork [8] to generate state-dependent parameters. However, such neural-network-based credit assignment functions cannot effectively incorporate prior knowledge, which may result in slower convergence and reduced interpretability.
@@ 4.1 Training-Free Credit Assignment Function
To address these issues, we propose a training-free credit assignment function (TFCAF), which serves as the core module of QLLM. This module is generated by LLMs based on a series of task-related prompts, with correctness ensured through structured code verification. As illustrated in Equation (2), the global Q-value function $Q_{\text{tot}}$ can theoretically be decomposed into a weighted sum of individual local Q-values with an additional bias term. Motivated by this decomposition paradigm, a similar formulation is adopted in the design of TFCAF. Its mathematical expression is as follows:
$Q_{\text{tot}}\left(s,\boldsymbol{a}\right)=\sum_{i=1}^{n}{f_{w}^{i}}\left(s\right)Q_{i}\left(\tau^{i},a^{i}\right)+f_{b}\left(s\right).$ (3)
The coefficient $f_{w}^{i}\left(s\right)$ is a state-dependent weight, and $f_{b}\left(s\right)$ is a bias term that also depends on the global state. Both are key components of the TFCAF, which are generated by LLMs. A detailed theoretical proof justifying this decomposition formulation is provided in Appendix 0.D. In the following, we provide a detailed explanation of how the TFCAF is generated, as well as how it facilitates credit assignment in multi-agent systems.
4.2 Coder-Evaluator Framework
Our framework employs two distinct LLM roles: the coder LLM $M_{coder}$ for generating candidate TFCAFs and the evaluator LLM $M_{evaluator}$ for critical selection and validation, ensuring the robustness of the generated functions with minimal human involvement.
@@ 4.2 Coder-Evaluator Framework
The interaction between the two models is driven by two types of prompts. Specifically, Task prompts ($P_{task}$) provides environment-specific details, while Role prompts ($P_{role}$) define the fixed responsibilities and operational roles of each LLM within the framework. These role prompts are further categorized into coder prompt and evaluator prompt, which remain unchanged across different RL tasks and correspond to the Coder LLM and the Evaluator LLM, respectively. This separation between task-specific and role-specific prompts enables QLLM to adapt to new cooperative tasks with minimal prompt modification, without requiring extensive manual tuning or subjective rewriting. Detailed prompt templates used in the framework are provided in Appendix 0.A.
@@
Initial Generation. At the initial stage, the Coder LLM $M_{\text{coder}}$ generates a set of candidate credit assignment functions conditioned on the provided prompts. Specifically, given the role-specific prompts $P_{\text{role}}$ and the task-specific prompts $P_{\text{task}}$, $M_{\text{coder}}$ produces $K$ candidate functions:
$\phi_{1},\phi_{2},\dots,\phi_{K}\ \Leftarrow\ M_{\text{coder}}\big(P_{\text{role}},P_{\text{task}}\big),$ (4)
where each $\phi_{k}$ represents a candidate TFCAF. Each generated function explicitly specifies how to compute the state-dependent weight functions $\{f_{w}^{i}(s)\}_{i=1}^{n}$ and the bias term $f_{b}(s)$ in Equation (3), typically in the form of executable code or programmatic expressions. Consequently, each $\phi_{k}$ defines a complete mapping from the global state $s$ to the mixing coefficients used for credit assignment, without introducing additional learnable parameters.
Error Detection and Correction. Due to the inherent tendency of LLMs to produce hallucinated or imperfect code, the generated candidate functions may contain syntax errors, dimensional mismatches, or invalid operations. To ensure executability, each candidate TFCAF is first compiled and executed once using a set of available input states and corresponding local action-value estimates. If an execution error occurs, the corresponding error message returned by the compiler or runtime environment is collected and provided as structured feedback to the Coder LLM $M_{\text{coder}}$. Conditioned on this feedback, $M_{\text{coder}}$ regenerates a corrected version of the faulty TFCAF:
$\phi_{\text{correct}}\ \Leftarrow\ M_{\text{coder}}\big(P_{\text{role}},\,P_{\text{task}},\,\text{error}\big).$ (5)
This procedure continues until all $K$ candidate TFCAFs execute successfully, ensuring basic syntactic and execution-level correctness before they are passed to subsequent evaluation stages.
Evaluation of TFCAF. After execution-level validation, the Evaluator LLM $M_{\text{evaluator}}$ evaluates the candidate TFCAFs to identify the most suitable one for credit assignment. Given the validated candidate set $\Phi=\{\phi_{1},\phi_{2},\dots,\phi_{K}\}$, $M_{\text{evaluator}}$ focuses on whether a TFCAF provides a coherent and task-aligned mapping from the global state to agent-wise credit weights, rather than relying on empirical performance metrics.
The evaluation and final choice are fully guided by structured role and task prompts and are performed automatically by $M_{\text{evaluator}}$, without manual ranking or subjective human intervention:
$\phi_{\text{choice}}\ \Leftarrow\ M_{\text{evaluator}}\big(P_{\text{role}},\,P_{\text{task}},\,\Phi\big),$ (6)
where the selected TFCAF is subsequently adopted as the credit assignment function in QLLM. Subsequently, $\phi_{\text{choice}}$ is fed back into $M_{\text{coder}}$ to serve as a reference for the next cycle of generation, and QLLM iteratively repeats this process for a total of $T$ rounds to ensure continuous logic refinement.
4.3 Training Procedure
@@
where $f_{\text{TFCAF}}$ is a parameterized function in which the weights are allowed to take arbitrary values, while the biases are also unconstrained. $f_{\text{TFCAF}}$ receives the individual action-value functions $Q_{i}$ as input and utilizes the global state $s$ to generate the corresponding weights $f_{w}^{i}$ and bias $f_{b}$ through the coder-evaluator framework, respectively, both of which are modeled as nonlinear functions of the global state. Notably, all parameters in the TFCAF are generated by LLMs and remain fixed during training. The global Q-value is then computed as shown in Equation (3). The training objective is to minimize the temporal-difference (TD) loss between the predicted total Q-value and the target Q-value:
@@
where the target value $y$ follows the Bellman equation:
@@
where $\theta^{-}$ denotes the parameters of the target agent network.
@@
5 Experiments
To comprehensively evaluate the efficacy and versatility of the proposed QLLM framework, we design and conduct experiments focused on the following objectives: (1) assessing the performance of QLLM across a diverse range of cooperative multi-agent environments; (2) validating the generality and compatibility of QLLM by integrating it with various established value decomposition algorithms; (3) evaluating its robustness to high-dimensional global state spaces by systematically increasing the number of agents and state dimensionality; and (4) investigating the interpretability of TFCAF to provide insights into the generated credit assignment logic. Furthermore, additional ablation studies and detailed comparisons of the number of learnable parameters across different algorithms are provided in Appendix 0.B.
5.1 Experimental Setups
@@
We tested QLLM’s performance on four common MARL benchmarks and compared it with several well-known baseline credit assignment algorithms. The primary LLM used in our method is the DeepSeek-R1 [7] inference model. For most experiments, the reported values represent the mean over five independent training runs, with 90% confidence intervals indicated by shaded areas. Notably, for SMAC scenarios, the solid lines represent the median test win rate, while the shaded regions denote the 25th-75th percentile interquartile range. Our experimental environments include: (1) Level-Based Foraging (LBF) [30], for testing in sparse reward settings; (2) Google Research Football (GRF) [16], to assess performance in complex, realistic environments; (3) Multi-agent Particle Environments (MPE) [23], for evaluating performance in high-dimensional state spaces; and (4) StarCraft Multi-Agent Challenge (SMAC) [36], to test performance in complex and highly cooperative multi-agent scenarios.
@@
We compare QLLM with seven baseline algorithms: QMIX [35], QPLEX [40], Qatten [43], RIIT [9], COMA [6], Weighted QMIX [34] and MASER [11]. A detailed introduction and implementation details for all baselines and environments are provided in Appendix 0.C.
5.2 The Superiority of QLLM
Figure 3 show the performance of various algorithms in LBF and GRF environments. It is evident that the QLLM algorithm outperforms the baseline in both the LBF environment and the GRF environment.
Figure 3: Performance comparison in LBF and GRF environments.
5.3 The Generality of QLLM
@@
QLLM is based on value decomposition methods, so it is compatible with most MARL algorithms that use mixing networks to compute global Q-value. To verify this compatibility, we extend our evaluation beyond QMIX, using the TFCAF generated by QLLM to replace the mixing networks in two other value decomposition algorithms: RIIT and MASER. We conducted experiments in the MPE environment, and the experimental results are shown in Figure 4. Our results demonstrate that QLLM can be integrated with various value decomposition algorithms to significantly enhance their performance.
Figure 4: Comparison of QLLM-enhanced variants with their original baselines in the MPE environments.
5.4 The Performance of QLLM in High Dimensional State Spaces
We evaluated the scalability of QLLM by increasing the number of agents to $N=15$ and $N=25$ in the simple-spread task. As shown in Figure 5, as state dimensionality grows, credit assignment becomes difficult for traditional value decomposition baselines, leading to significant performance degradation. In contrast, QLLM maintains high contribution assignment accuracy. This robustness stems from the LLM-generated nonlinear TFCAFs, which leverage well-defined task logic that is invariant to dimensionality, unlike traditional mixing network that struggle with high-dimensional parameter optimization.
Figure 5: The results of high-dimensional testing in simple-spread task.
5.5 The Performance on SMAC
We evaluated QLLM on two hard SMAC maps: 3s_vs_5z and 2c_vs_64zg. As illustrated in Figure 6, QLLM outperforms all baselines with significantly faster convergence. These results demonstrate that LLM-generated logic efficiently solves complex coordination tasks where traditional neural mixers typically struggle to optimize during the early stages of training.
Figure 6: Performance comparison on SMAC maps.
5.6 The Interpretability of TFCAF
To demonstrate the interpretability of QLLM, Listing 5.6 presents the complete generated TFCAF for the academy_3_vs_1_with_keeper scenario, while Figure 7 provides a visualization and a detailed explanation of its weight components. The generated logic explicitly leverages physical factors, such as ball possession and goal proximity, to attribute credit. Unlike traditional neural mixing networks that operate as black boxes, TFCAF offers a transparent and human-readable mechanism, clearly revealing the underlying tactical rationale for credit assignment in complex environments.
{listing}
[htbp] TFCAF for the academy_3_vs_1_with_keeper scenario in GRF.
def QLLMNetwork(agents_q, global_state): # Extract environment informations ball_x, ball_y = global_state[:, 88], global_state[:, 89] possession, holder_my_team=global_state[:,94:97],global_state[:,97:101] ball_goal_dist = torch.sqrt((ball_x-1.0)**2 + ball_y**2) my_team_has_ball = torch.argmax(possession, dim=1) == 0 in_push_region = (ball_goal_dist>0.19) & (ball_goal_dist<0.99) & \ my_team_has_ball agent_x, agent_y = global_state[:,[0,2,4,6]],global_state[:,[1,3,5,7]] agent_goal_dist = torch.sqrt((agent_x-1.0)**2 + agent_y**2) agent_ball_dist = torch.sqrt((agent_x-ball_x.unsqueeze(1))**2 + (agent_y-ball_y.unsqueeze(1))**2) # Calculate weights components holder_weights = (1.0-ball_goal_dist).unsqueeze(1)*holder_my_team * \ in_push_region.unsqueeze(1)*10.0 non_push_weights = holder_my_team*my_team_has_ball.unsqueeze(1) * \ (~in_push_region).unsqueeze(1) support_weights=(1.0/(1.0+agent_goal_dist))*(~holder_my_team.bool())* \ my_team_has_ball.unsqueeze(1) defense_weights = (1.0/(1.0+agent_ball_dist)) * \ (~my_team_has_ball).unsqueeze(1) weights = torch.softmax(holder_weights + non_push_weights + support_weights + defense_weights, dim=1) return (agents_q * weights).sum(dim=1, keepdim=True)
Figure 7: Visualization and Tactical Explanation of Weight Components in the TFCAF for academy_3_vs_1_with_keeper scenario.
6 Conclusion